# Deep Correlated Prompting for Visual Recognition with Missing Modalities

### v2

@@ 4.2 Experimental Results
Figure 3: Comparison of our final model (Ours) with (1) baseline, which directly drops the features when a modality is missing; (2) Ours (A), which only equips the correlated prompts; (3) Ours (B), which equips both the correlated prompts and the dynamic prompts. The experiments are conducted on the MM-IMDb dataset [2] across different missing rates (0–100%) upon three different missing-modality scenarios (missing-both, missing-image and missing-text).
@@
Ablation study. We test the configurations for our proposed three prompts in Tab. 4.2, Tab. 4.2 and Tab. 4.2, respectively. In the upper part of Tab. 4.2, we first explore how to leverage the correlations between prompts of different layers. "No projection" denotes using independent prompts for each layer and "Fc" denotes using a fully connected layer for projection. It’s observed that compared to using independent prompts, either using a fc or a MLP offers a notable performance boost. The MLP gives better performance for its non-linearity and we use it by default. We then explore the prompt depth $J$ (how many layers to insert prompts) in the middle of Tab. 4.2. We notice that the performance reaches a peak when $J$=6, and either decreasing it or increasing it would degrade the performance. Finally, we verify the efficacy of incorporating multimodal prompts for instruction in the bottom part of Tab. 4.2. It’s observed that introducing bi-modal prompts to generate the prompts of the current layer gives better performance, which shows that different modalities can offer complementary information.
@@
Figure 4: Ablations on the generalizability to different testing scenarios across various missing rates on the MM-IMDb dataset [2]. (a) All models are trained on missing-both cases, and evaluated on missing-both cases with different missing rates. (b) Models are trained on missing-both or missing-image cases, and evaluated on missing-image cases with different missing rates. (c) Models are trained on missing-both or missing-text cases, and evaluated on missing-text cases with different missing rates.
@@
Datasets Missing rate $\eta$ Training Testing CoOp MMP MaPLe DePT Ours
Image Text Image Text
MM-IMDb (F1-Macro) $50\%$ $100\%$ $50\%$ $100\%$ $50\%$ 51.23 52.07 52.76 53.87 55.23
$50\%$ $100\%$ $50\%$ $100\%$ 53.04 54.52 55.26 56.04 57.32
$75\%$ $75\%$ $75\%$ $75\%$ 51.46 52.12 52.87 54.02 55.45
$70\%$ $100\%$ $30\%$ $100\%$ $30\%$ 47.26 48.23 48.75 49.87 51.35
$30\%$ $100\%$ $30\%$ $100\%$ 52.32 53.21 53.98 55.04 56.21
$65\%$ $65\%$ $65\%$ $65\%$ 50.22 51.34 52.31 53.17 54.24
$90\%$ $100\%$ $10\%$ $100\%$ $10\%$ 47.86 37.36 50.12 50.98 52.36
$10\%$ $100\%$ $10\%$ $100\%$ 51.65 52.36 53.14 54.12 55.42
$55\%$ $55\%$ $55\%$ $55\%$ 47.44 48.04 48.82 49.98 51.26
Food101 (Accuracy) $50\%$ $100\%$ $50\%$ $100\%$ $50\%$ 77.36 78.24 79.87 80.24 82.33
$50\%$ $100\%$ $50\%$ $100\%$ 86.98 87.12 87.48 87.85 89.23
$75\%$ $75\%$ $75\%$ $75\%$ 81.76 81.98 82.58 83.26 85.25
$70\%$ $100\%$ $30\%$ $100\%$ $30\%$ 76.65 76.74 76.87 76.87 79.18
$30\%$ $100\%$ $30\%$ $100\%$ 85.21 86.12 86.36 86.52 87.53
$65\%$ $65\%$ $65\%$ $65\%$ 79.14 79.56 80.06 81.85 82.38
$90\%$ $100\%$ $10\%$ $100\%$ $10\%$ 72.65 73.74 73.25 74.22 75.54
$10\%$ $100\%$ $10\%$ $100\%$ 82.16 82.78 83.42 84.02 86.26
$55\%$ $55\%$ $55\%$ $55\%$ 77.36 77.78 78.26 78.66 80.39
Hateful Memes (AUROC) $50\%$ $100\%$ $50\%$ $100\%$ $50\%$ 60.56 60.31 60.87 61.87 62.32
$50\%$ $100\%$ $50\%$ $100\%$ 62.41 62.35 63.13 63.88 64.46
$75\%$ $75\%$ $75\%$ $75\%$ 64.87 65.84 65.46 65.86 66.02
$70\%$ $100\%$ $30\%$ $100\%$ $30\%$ 60.74 61.12 61.26 61.56 62.82
$30\%$ $100\%$ $30\%$ $100\%$ 62.74 63.24 63.14 63.48 64.12
$65\%$ $65\%$ $65\%$ $65\%$ 64.82 65.04 65.23 65.48 66.08
$90\%$ $100\%$ $10\%$ $100\%$ $10\%$ 60.03 57.21 60.74 61.14 62.08
$10\%$ $100\%$ $10\%$ $100\%$ 61.46 61.52 61.87 62.42 63.87
$55\%$ $55\%$ $55\%$ $55\%$ 64.32 63.34 64.85 65.37 66.78
@@
Table 5: Performance v.s. MMP when owning comparable parameters
@@ 48. Answer: [Yes]
  49. Justification: They are discussed in the appendix.

### v4

@@ 4.2 Experimental Results
Figure 3: Comparison of our final model (Ours) with (1) baseline, which directly drops the features when a modality is missing; (2) Ours (A), which only equips the correlated prompts; (3) Ours (B), which equips both the correlated prompts and the dynamic prompts. The experiments are conducted on the val set of MM-IMDb dataset [2] across different missing rates (0–100%) upon three different missing-modality scenarios (missing-both, missing-image and missing-text).
@@
Ablation study. We test the configurations for our proposed three prompts on the val set of MM-IMDb dataset in Tab. 4.2, Tab. 4.2 and Tab. 4.2, respectively. In the upper part of Tab. 4.2, we first explore how to leverage the correlations between prompts of different layers. "No projection" denotes using independent prompts for each layer and "Fc" denotes using a fully connected layer for projection. It’s observed that compared to using independent prompts, either using a fc or a MLP offers a notable performance boost. The MLP gives better performance for its non-linearity and we use it by default. We then explore the prompt depth $J$ (how many layers to insert prompts) in the middle of Tab. 4.2. We notice that the performance reaches a peak when $J$=6, and either decreasing it or increasing it would degrade the performance. Finally, we verify the efficacy of incorporating multimodal prompts for instruction in the bottom part of Tab. 4.2. It’s observed that introducing bi-modal prompts to generate the prompts of the current layer gives better performance, which shows that different modalities can offer complementary information.
@@
Figure 4: Ablations on the generalizability to different testing scenarios across various missing rates on the val set of MM-IMDb dataset [2]. (a) All models are trained on missing-both cases, and evaluated on missing-both cases with different missing rates. (b) Models are trained on missing-both or missing-image cases, and evaluated on missing-image cases with different missing rates. (c) Models are trained on missing-both or missing-text cases, and evaluated on missing-text cases with different missing rates.
@@
Datasets Missing rate $\eta$ Train/Test Validation set Testing set
Image Text CoOp MMP MaPLe DePT Ours CoOp MMP MaPLe DePT Ours
MM-IMDb (F1-Macro) $50\%$ $100\%$ $50\%$ 51.23 52.07 52.76 53.87 55.23 48.06 48.88 49.58 50.64 52.13
$50\%$ $100\%$ 53.04 54.52 55.26 56.04 57.32 49.89 51.46 52.32 52.78 54.32
$75\%$ $75\%$ 51.46 52.12 52.87 54.02 55.45 48.37 49.32 49.56 50.87 52.32
$70\%$ $100\%$ $30\%$ 47.26 48.23 48.75 49.87 51.35 44.13 45.64 45.52 46.38 48.52
$30\%$ $100\%$ 52.32 53.21 53.98 55.04 56.21 48.82 50.52 50.64 52.13 53.14
$65\%$ $65\%$ 50.22 51.34 52.31 53.17 54.24 46.84 48.12 49.16 50.32 51.42
$90\%$ $100\%$ $10\%$ 47.86 48.84 50.12 50.98 52.36 44.76 45.32 46.84 47.56 49.26
$10\%$ $100\%$ 51.65 52.36 53.14 54.12 55.42 48.32 49.12 50.13 50.88 52.22
$55\%$ $55\%$ 47.44 48.04 48.82 49.98 51.26 44.12 44.87 45.12 46.54 48.04
Food101 (Accuracy) $50\%$ $100\%$ $50\%$ 77.36 78.24 79.87 80.24 82.33 77.45 77.89 79.64 80.16 82.11
$50\%$ $100\%$ 86.98 87.12 87.48 87.85 89.23 87.02 87.16 87.35 82.14 89.12
$75\%$ $75\%$ 81.76 81.98 82.58 83.26 85.25 81.24 81.72 82.34 83.12 85.24
$70\%$ $100\%$ $30\%$ 76.65 76.74 76.87 76.87 79.18 76.34 76.52 77.02 77.34 78.87
$30\%$ $100\%$ 85.21 86.12 86.36 86.52 87.53 84.78 85.64 85.89 86.12 87.32
$65\%$ $65\%$ 79.14 79.56 80.06 81.85 82.38 78.87 79.12 79.84 81.46 81.87
$90\%$ $100\%$ $10\%$ 72.65 73.74 73.25 74.22 75.54 71.87 73.14 73.46 74.12 75.26
$10\%$ $100\%$ 82.16 82.78 83.42 84.02 86.26 81.67 82.14 83.12 83.56 85.78
$55\%$ $55\%$ 77.36 77.78 78.26 78.66 80.39 76.46 76.58 77.85 78.12 79.87
Hateful Memes (AUROC) $50\%$ $100\%$ $50\%$ 58.32 58.56 58.78 59.31 60.24 60.56 60.31 60.87 61.87 62.32
$50\%$ $100\%$ 60.34 61.12 61.34 61.78 62.34 62.41 62.35 63.13 63.88 64.46
$75\%$ $75\%$ 62.34 62.87 63.14 63.24 63.78 64.87 65.84 65.46 65.86 66.02
$70\%$ $100\%$ $30\%$ 58.54 59.02 59.36 60.02 60.56 60.74 61.12 61.26 61.56 62.82
$30\%$ $100\%$ 60.12 60.78 61.32 61.54 62.32 62.74 63.24 63.14 63.48 64.12
$65\%$ $65\%$ 62.34 62.56 63.12 63.32 63.78 64.82 65.04 65.23 65.48 66.08
$90\%$ $100\%$ $10\%$ 58.02 57.34 58.32 59.02 60.34 60.03 57.21 60.74 61.14 62.08
$10\%$ $100\%$ 59.02 59.32 60.21 60.56 61.34 61.46 61.52 61.87 62.42 63.87
$55\%$ $55\%$ 62.32 62.56 63.24 63.78 64.34 64.32 63.34 64.85 65.37 66.78
@@
Table 5: Performance v.s. MMP when owning comparable parameters on the val set of MM-IMDb [2].
@@ 48. Answer: [Yes]
  49. Justification: They are discussed at the end of the paper.